OpenStudy (kimberly_pr):

Help with 3 questions??

OpenStudy (kimberly_pr):

What are the explicit equation and domain for a geometric sequence with a first term of 2 and a second term of −8? an = 2(−8)n − 1; all integers where n ≥ 1 an = 2(−8)n − 1; all integers where n ≥ 0 an = 2(−4)n − 1; all integers where n ≥ 0 an = 2(−4)n − 1; all integers where n ≥ 1

OpenStudy (kimberly_pr):

The function f(x) = 16(2)x represents the growth of a bee population every year in a remote swamp. Jennifer wants to manipulate the formula to an equivalent form that calculates two times a year, not just once a year. Which function is correct for Jennifer's purpose, and what is the new growth rate? f(x) = 16(2)x; growth rate 200% f(x) = 16(2)2x; growth rate 8% f(x) = 16(1.41)x; growth rate 8% f(x) = 16(1.41)2x; growth rate 41%

OpenStudy (kimberly_pr):

What is the solution of log2x + 6 144 = 2? x = −6 x = −3 x = 3 x = 6

OpenStudy (kimberly_pr):

Nevermind, I guessed and got 2 wrong and 1 right. I guess I'm failing Algebra 2

OpenStudy (kimberly_pr):

Jk, I got an 80 on the assignment overall

OpenStudy (agent0smith):

lol one question at a time next time.